Two trains T1 & T2 travelling in opposite directions,on parallel tracks,are just beginning to pass one another. Train T1 is 72 metres longand is travelling at 108 kilometres par hour (108km/h. T2 is 78 metres long and is travelling at 72 kilometres per hour (72km/h). Please assist in getting the time in seconds the trains take to completely pass each other
Answer by mananth(16949)

Train T1 is 72 metres long
and is travelling at (108km/h.
T2 is 78 metres long and is travelling (72km/h).
T1 & T2 are travelling in opposite directions
so add up speed = 108+72=180 km/h
Distance = sum of lengths of trains (72+78)=150 m= 0.15km
time = d/r
t=0.15/180
t=0.00083 hours
1 hour = 3600 seconds
0.000833*3600= 3 seconds
Time to pass= 3 seconds
m.ananth@hotmail.ca